OpenStudy (anonymous):

find exact value of sin(x+pi/6) if sinx=-12/13 and tan<0

OpenStudy (freckles):

Expand sin(x+pi/6) using sum identity for sin You will also need to find cos(x) Recall \[\cos^2(x)=1-\sin^2(x) \\cos(x)=\pm \sqrt{1-\sin^2(x)}\] You will need to determine if cos(x)>0 of if cos(x)<0. You can do find knowing that sin is negative and tangent is negative, forcing cosine to be?

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

So would I use the equation sin(pi/3+pi/2)=sin(pi/3)cos(pi/2)+cos(pi/3)sin(pi/2)?

OpenStudy (anonymous):

or would it be sin(x+pi/6)=sin(-12/13)cos(11pi/6)+cos(-12/13)sin(11pi/6)?

OpenStudy (freckles):

sin(x+pi/6) =sin(x)cos(pi/6)+sin(pi/6)cos(x)

OpenStudy (freckles):

you are not given x is -12/13 you are given sin(x) is -12/13

OpenStudy (freckles):

I gave you an equation to find cos(x)

OpenStudy (anonymous):

is this correct? sin(-12/13)cos(sqrt3/2)+cos(5/13)sin(1/2)

OpenStudy (freckles):

Don't replace x with -12/13 replace sin(x) with -12/13 and I don't think x is 5/13

OpenStudy (freckles):

cos(pi/6) is sqrt(3)/2 not cos(sqrt(3)/2)

OpenStudy (freckles):

sin(pi/6)=1/2 not sin(1/2)

OpenStudy (freckles):

\[\sin(x)=\frac{-12}{13} \\ \cos(x)=\sqrt{1-(\frac{-12}{13})^2}=\sqrt{1-\frac{144}{169}}=\sqrt{\frac{169-144}{169}}=\sqrt{\frac{25}{169}}=\frac{5}{13}\]

OpenStudy (freckles):

So this tells us cos(x) is 5/13 not that x is 5/13

OpenStudy (freckles):

\[\sin(x+\frac{\pi}{6})=\sin(x)\cos(\frac{\pi}{6})+\cos(x)\sin(\frac{\pi}{6}) \\ =\frac{-12}{13} \cdot \frac{\sqrt{3}}{2}+\frac{5}{13} \cdot \frac{1}{2}\]

OpenStudy (freckles):

Finish simplifying that.

OpenStudy (anonymous):

That is what I did and I came up with -12sqrt3+5/26

OpenStudy (freckles):

Ok well you were replacing x with -12/13 and 5/13 even though it is already weird to say x is taking on two different values at the same time So that was very concerning...

OpenStudy (freckles):

And the other weird thing is cos(pi/6) =sqrt(3)/2 not cos(sqrt(3)/2)) and sin(pi/6)=1/2 not sin(1/2)

OpenStudy (anonymous):

I think I was just typing it differently which made it look confusing?

OpenStudy (freckles):

Not just differently but incorrectly.

OpenStudy (freckles):

Don't let your teacher see that way. :p

OpenStudy (anonymous):

ok thank you